If the supply corn is Qn=b+0.2p and the supply by the rest of the world is Qr=e+0.8p, what is the world supply? What is the world inverse supply?


1
Expert's answer
2022-03-23T18:34:09-0400

Qn=b+0.2p and Qr=e+0.8p

find average gradient:

"\frac{0.2+0.8}{2}" =0.5

therefore the world supply will be

Qw=e+0.5p

inverse supply

Qw=e+0.5p

0.5p=q-e

p=2q-2e